OpenStudy (anonymous):

find the inverse laplace transform of F(s)= (s+5)/((s^2+9)^2)

OpenStudy (zzr0ck3r):

first I would do partial fractions to expand

OpenStudy (zzr0ck3r):

so (s+5)/(s^2+3^2)^2 = s/(s^2 + 3^2)^2 + 5/(s^2+3^2)^2

OpenStudy (zzr0ck3r):

on tthe left you want it to be the form os t*sin(wt)

OpenStudy (zzr0ck3r):

so s/(s^2+3^2)^2 = ((2s(3))/6(s^2+3^2)^2

OpenStudy (zzr0ck3r):

= t*sin(3t)

OpenStudy (zzr0ck3r):

err = (1/6)(t*sin(w*t))

OpenStudy (zzr0ck3r):

you with me so far?

OpenStudy (anonymous):

sure ..continue please

OpenStudy (anonymous):

Hellooooo ??